PDCCODE—Documented TEX Macros

Version 1.6, last changed pdc 1995–03–30

P. Damian Cugley

1. Introduction

PDCCODE is a simpler and less featureful system than the LaTEX 2e Docstrip package, but uses the same file suffix ‘dtx’. Lines between \code and \endcode markers are written to a file (set with the \codefile command), as well as being printed ‘verbatim’ in document. Thus one TEX pass generates both the macro file(s) and a dvi file.

This document—generated from pdccode.dtx—is a dtx file for PDCCODE itself, the file of macros used by my dtx files. It cannot be used to generate pdccode.tex directly (TEX might choke halfway through overwriting pdccode.tex with a new version, leaving no pdccode.tex to generate pdccode.tex with). Instead it produces its copy as ‘tmp.tex’, and ‘tmp.tex’ can be compared to the supplied pdccode.tex.

1.1. How to read this document

The dtx files are programs (in TEX’s macro language) that generate new programs (the macro files that will be used in documents). The macros described in this file are used when generating macro files. So, for example, the \rcs and \fileversion macros used to extract identification information from the dtx file are not defined in any of the generated macro files.

1.2. Background

For some time I have been using a collection of macro files to typeset various types of document with TEX. After using them in Malvern* releases I have discovered that they had been merged with the UnixTEX macros directory. Originally I had never intended that these files be used by anyone but me—they were full of hacks and messy code and definitions I don’t use anymore. Since then I thought of tidying up the macros enough that I’m not embarassed by them, but the task of doing this and documenting their use was daunting.

After installing LaTEX 2e on my new computer, I decided to have a go at adapting the idea of ‘documented TEX macros’ files to plain TEX. This way I could clean up the macros and document them in a systematic fashion.

The LaTEX Docstrip/Doc-option approach evolved through several different models and so is a complex format. Docstrip (a program written in TEX’s macro language) was used to strip comments from doc files making sty files, and the LaTEX Doc-option separately extracted LaTEX code from comments in doc files to make printed documentation. Both Doc and Docstrip required the user to type in file names whenever they ran—which made them useless for batch processing. So separate driver files (with suffixes drv and ins) were added to distributions. Finally, the doc files, which used to be usable as sty files (just a little slower because of the comments) were have been changed to be documents instead (making drv files redundant, once the ltxdoc.cfg configuration file was added to the conventions) and given a new suffix dtx.

The system I use—called PDCCODE—is simpler and faster, but omits indexes, special change-log commands, and all the features that come with LaTEX formatting. It

– PDCCODE—Documented TEX Macros 1.6 1995–03–30

* Malvern is a sanserif font family implemented in METAFONT. The current version is release 1.2. doesn’t have Docstrip’s options feature, which allows several largely-similar files to be generated from the same source.† It doesn’t print statistics at the end, or do checksums, or even print comforting lines of asterisks and percent signs in the log file.

My dtx files use a single macro file pdccode.tex which defines a ‘code mode’ where lines are written to the macro file as well as appearing in the printed documentation. One pass generates both the macro file and the documentation in dvi form. The pdccode.tex file also includes enough layout commands to typeset the comments.

2. Summary of using PDCCODE

These commands are explained in more detail below.

2.1. A skeleton code file A code file using pdccode.tex will look something like this:

% name -- one line description

%%%@TeX-document-file {

header comments for the document

%%%}
\relax
\input pdccode
\document

\rcs$ RCS identification string $\endrcs

\codefile{ file name }

\author{ author name } \title{ title for the document }

\section{Introduction}

introduction

\section{File identification}
\code
  \|\% file name \fileversion~\filedate~-- one line desc
  |
  |%%%@TeX-definition-file {

header comments for macro file

  |%%%}
  |
  \|\\message\{\fileversion~<\fileauthor~\filedate>\}
  |
\endcode

definitions and commentary

\endcodefile
\enddocument
\bye

2.2. Commands in code sections Code sections start with \code and end with \endcode. Before the first \code you must do \codefile{⟨file name⟩} to specify the file.

Vertical mode material may be used between lines (e.g., \smallbreaks). Unexpanded code lines are introduced with |, expanded lines with \|. In expanded lines, use \{, \}, \\, \%, \~ for those characters, ~ for a single space.

† I do have an experimental version pdccode2.tex which does this but I am not particularly happy with it yet.

2.3. Commands in commentary

Most of the commentary will be typeset using the conventions of plain TEX. \author{⟨text⟩} (must precede \title) \title{⟨heading⟩}

  \section{⟨heading⟩}
  \subsec{⟨heading⟩}
  \noindenttrue
  \footnote

\defverbatim⟨one-character control sequence⟩

$⟨number⟩\mm$ $⟨number⟩\em$ $⟨number⟩\pt$ $⟨number⟩\inch$ \<⟨name⟩> \cs{⟨name⟩} \arg{⟨argument name⟩}

  \TeX \LaTeX \LaTeX~2e
  \notepar ⟨paragraph⟩ \endnotepar

3. File identification

Nowadays, macro files start with some comments identifying the file.

 1 % pdccode.tex 1.6 1995/03/30 -- self-documenting TeX macros

2

 3 %%%@TeX-definition-file {
 4 %%% filename       = "$texmf/tex/plain/pdcmac/pdccode.tex",
 5 %%% version        = "1.6",
 6 %%% date           = "1995/03/30",
 7 %%% package        = "pdcmac 1.0",
 8 %%% author         = "P. Damian Cugley",
 9 %%% email          = "damian.cugley@comlab.ox.ac.uk",
10 %%% address        = "Oxford University Computing Laboratory,
11 %%%                   Parks Road, Oxford  OX1 3QD, UK",
12 %%% codetable      = "USASCII",
13 %%% keywords       = "TeX, plain TeX, macros",
14 %%% supported      = "Maybe",
15 %%% abstract       = "Macros for documented TeX macros (DTX files).
16 %%%                   This file was generated by running
17 %%%                   plain TeX on pdccode.dtx",
18 %%% copyright      = "Copyright (c) 1991-1995 P. Damian Cugley",
19 %%% copying        = "DO NOT DISTRIBUTE THIS FILE.
20 %%%                   Distribute pdccode.dtx only as part of the
21 %%%                   package it came in.",
22 %%% dependencies   = "",
23 %%% }

24

25 \message{1.6 <pdc 1995/03/30>}

26

We start by making a nickname for a scratch register. This is \toks@ in plain TEX; this definition gives it a name I can use without catcode hackery.

27 \toksdef\toksa=0

4. RCS Support

RCS (Revision Control System) is a free source code control system which is available from the Free Software Foundation. It inserts keyword-value pairs enclosed in dollar signs into the source code. This command extracts the file’s version, date, author and locker (if there is one). It is used like this:

\rcs$Id:␣⟨name⟩,v␣⟨version⟩␣⟨date⟩␣⟨ignored⟩ ␣⟨author⟩␣⟨ignored⟩␣[⟨locker⟩␣]$\endrcs The ⟨locker⟩ field is non-empty iff the file is ‘locked’ (meaning it has been edited and the changes are not checked in yet). The \filename macro is defined, even though it will almost certainly be identical to \jobname.

28 \def\rcs#1: #2,v #3 #4 #5 #6 #7 #8$#9\endrcs{
29  \def\filename{#2}
30  \def\fileversion{#3}
31  \def\filedate{#4}
32  \def\fileauthor{#6}
33  \def\filelocker{#8}

34 }

5. Verbatim text

Verbatim text is text that is printed identically to the characters in the manuscript file. In the code sections, a vertical bar is used to make lines for the code file. In the documentation, the vertical bar is used as a self-matching verbatim mode character.

These definitions should be the same as the corresponding code in pdcfmt2.dtx.

5.1. Lists of active characters The verbatim text macros need to know if any characters have been made special apart from the usual ones. This command tells the verbatim macros that #1 is an active char; in addition to setting its catcode, this makes sure that the character will be re-catcoded appropriately when in verbatim mode. #1 should be a something that will work after ‘` ’; usually this means a one-character control sequence.

35 \def\declareactivechar#1{%
36  \toksa\expandafter{\verbatimplains\do#1}%
37  \edef\verbatimplains{\the\toksa }%
38  \catcode` #1\active

39 }

Two list macros in the style of \dospecials are defined. The first is for characters that must be set to catcode 12, such as ‘\’, ‘{’, ‘}’, ‘_’, ‘$’, ‘#’, ‘&’, and ‘%’. The other is for a few characters that are given catcode 13, such as ‘` ’, ‘´ ’, ‘~’, ‘^’ and SP.

40 \def\verbatimplains{\do\\\do\{\do\}\do\_\do\$\do\#\do\&\do\%}
41 \def\verbatimactives{\do\` \do\´ \do\~\do\^\do\ }
42 \chardef\other=12

5.2. Setting up verbatim mode This is a macro common to various verbatim text modes. The default spacing of whatever font is used, without any adjustments of spacefactor for punctuation. The characters ‘` ’, ‘´ ’, ‘~’, and ‘^’ produce slightly lowered glyphs.

43 \newtoks\everyverbatim
44 \bgroup \catcode` \^=13 \catcode` \´ =13 \catcode` \` =13 \toksa={\egroup
45  \def\setupverbatim{%
46  \frenchspacing
47  \spaceskip0pt \xspaceskip0pt % use spacing of font
48  \def\do##1{\catcode\lq##112 }\verbatimplains
49  \def\do##1{\catcode\lq##1\active }\verbatimactives
50  \let` \ttlq \let´ \ttrq
51  \let~\tttilde \let^\ttcircum
52  \the\everyverbatim

53 }

54 }\the\toksa

Here are the definitions for the munged glyphs. These are only necessary in CM-Typewriter-like fonts; using something like Courier, these definitions could probably be omitted.

The names ‘rq’ and ‘lq’ are based on the use of \lq and \rq in plain TEX. An alternative definition for \ttrq would be \chardef\ttrq=13, producing ‘'’.

55 \def\ttlq{\lower0.125ex \hbox{\char18 }}

 56 \def\ttrq{\lower0.125ex \hbox{\char19 }}
 57 \def\tttilde{\lower0.5ex \hbox{\char` \~ }}
 58 \def\ttcircum{\lower0.5ex \hbox{\char` \^}}

This makes the difference between ` ´ ^~ and `'^~.*

5.3. Defining a self-matching character for verbatim text

A macro for defining a self-matching verbatim character (I got the idea of a macro for declaring verbatim characters from the LaTEX 2e documentation macros). #1 must be a one-character control sequence. If it isn’t active already, then \declareactivechar is applied to it first.

 59 \def\defverbatim#1{%
 60  \ifcat\noexpand#1\noexpand~\else \declareactivechar#1 \fi
 61  \begingroup \uccode` \~=` #1 \uppercase{\toksa={\endgroup
 62  \def~{%
 63  \leavevmode
 64  \begingroup \tt \setupverbatim
 65  \catcode` #1\active \let~\endgroup
 66  }%
 67  }}\the\toksa

68 }

Note that there must be a font called \tt for it to switch to. (The catcode of #1 is set to 13 after calling \setupverbatim because #1 will be included in \verbatimplains.)

Finally, we define vertical bar as the conventional self-matching verbatim-mode character.

 69 \defverbatim\|
 70 \hyphenchar\tentt=-1
    Both \defverbatim and \defineactivechar have local scope.

6. Code sections

The markers \code … \endcode mark code to be written to the definition file. Each line starts with either | or \| and finishes with the end of the line in the manuscript file.

Lines that start with | are written out as-is; no macro expansion is allowed. This is how the vast majority of lines are done.

Lines starting \| are expanded. This is so that things like the current version ID (extracted from the RCS identification line or set with \def) may be included in the code file. Within such an expanded line, the following macros are defined:

\%  % per cent sign \{  { left brace \}  } right brace
\\  \ backslash  ~   ␣ space  \~  ~ tilde

(This list is probably incomplete.)

6.1. Setting the codefile name

Before the first code section, the \codefile{⟨file name⟩} command specifies the code file. The last code section should be followed by \endcodefile. Only one code file may be open at a time.

 71 \newwrite\CODEwrite

* It turns out that on some TEX systems, the CMTT font has been hacked to have an ASCII-style reverse-apostrohe-cum-grave-accent as character 96 and a matching apostrophe-cum-acute-accent as charcter 39, thereby correcting the defects that my definitions of \ttlq and \ttrq address. But such modified fonts are nonstandard—they should not be called CMTT—and should be replaced.

72 \def\codefile#1{
73  \ifeof\CODEwrite\else \endcodefile \fi
74  \def\thecodefile{#1} \openout\CODEwrite=#1

75 }

76 \def\endcodefile{
77  \message{Code written on \thecodefile.}
78  \closeout\CODEwrite

79 }

6.2. Code sections

Switching to ‘code mode’ is easy, because most of the work is really done by the | and \| commands.

80 \def\code{
81  \smallskip
82  \begingroup
83  \let|\CODEline \let\|\CODElinex
84  \let\{=\CODElb \let\}=\CODErb \let\\=\CODEbslash
85  \def~{ } \let\~\CODEtilde
86  \the\everycode
87  \obeylines

88 }

89 \def\endcode{
90  \smallbreak
91  \endgroup

92 }

6.3. Code lines

This is the definition for unexpanded code lines.

93 \def\CODEline{%
94  \begingroup % matched by \CODElinei
95  \setupverbatim \catcode` \|\other
96  \CODElinei

97 }

In \CODElinei, We must redefine the characters (such as ` ) that are active within verbatim text to expand to their catcode-12 equivalents when they are written to the code file currectly.

Since one of these is ^, we temporarily give @ the meaning of ^so that we may safely change the catcode of ^.

98 \begingroup\catcode` \´ =\active \catcode` \` =\active
99  \catcode\lq\@=\catcode\lq\^\catcode\lq\^=\active
100  \catcode\lq\@@M=\active \toksa={\endgroup %
101  \def\CODElinei#1@@M{%
102  \begingroup\let` \lq \let´ \rq %
103  \let^\CODEcircum \let~ \CODEtilde%
104  \edef\tmp{#1}%
105  \write\CODEwrite\expandafter{\tmp}%
106  \endgroup %
107  \indent\the\everycodeline #1\par %
108  \endgroup % matches \CODEline

109 }%

110 }\the\toksa

The definition for \CODElinex (expanded code lines) is much simpler:

111 \begingroup\catcode` \^^M=13 \toksa={\endgroup %
112  \def\CODElinex#1^^M{%
113  \edef\tmp{\write\CODEwrite{#1}}\tmp %
114  {\indent\the\everycodeline #1\par}%

115 }%

116 } \the\toksa

6.4. Macros for making category-12 characters For characters that are active in verbatim mode (like ` ), or which must be specially generated in expanded code lines (such as braces generated with \{), we have a ‘\CODE-’ macro that generates an explicit character token with catcode 12.

117 \begingroup\catcode` \/=0 \catcode` \\=\other /toksa={/endgroup
118  /def/CODEbslash{\}
119 }/the/toksa
120 \begingroup\catcode` \^=\other \toksa={\endgroup
121  \def\CODEcircum{^}
122 }\the\toksa
123 \begingroup\catcode` \~=\other \toksa={\endgroup
124  \def\CODEtilde{~}
125 }\the\toksa
126 \begingroup\catcode` \%=\other \toksa={\endgroup
127  \def\%{%}
128 }\the\toksa
129 \begingroup\catcode` \~=\other \toksa={\endgroup
130  \def\CODEtilde{~}
131 }\the\toksa
132 \begingroup\catcode` \[=1 \catcode` \]=2
133  \catcode` \{=\other \catcode` \}=\other
134  \toksa=[\endgroup
135  \def\CODElb[{] \def\CODErb[}]
136 ]\the\toksa

6.5. Customization The token list register \everycode is expanded at the start of every code section, the token list register \everycodeline at the start of every code line. The default definitions select typewriter font for the code listings, with each line numbered.

137 \newtoks\everycode
138 \newtoks\everycodeline
139 \everycode{\tt \baselineskip=0.8\baselineskip \lineskiplimit=0pt }
140 \newcount\lineno
141 \everycodeline{\llap{$ \global\advance\lineno1
142  \scriptstyle\the\lineno $\enspace}%
143  \mark\expandafter{\the\lineno}}

7. Layout

The PDCCODE macros include enough basic formatting commands to make a simple document—since this file needs to be self-contained. The formatting is fairly primitive, so that pdccode.tex will not be excessively long.

The text is set in the default plain TEX fonts: 10-point Computer Modern Roman. Titles are in CM Sans Serif Bold Extended, headings in CMSS Demibold Condensed, and subheadings in CMSS Italic: all of these are standard fonts which should be available on all TEX systems.

The line width is wide enough that code lines do not need to be printed in a reduced font (which looks silly).

the first tokenm of a document is \relax and everything except the initial \input and final \bye in the document is enclosed in \document… \enddocument. This is so that documents using the PDCCODE conventions can be embedded within a another document which by redefining these tokens.

144 \let\document\relax \let\enddocument\relax

I have to set some TEX 3 parameters because we don’t seem to have definitions for them in the local version of plain.tex.

145 \lefthyphenmin=3 \righthyphenmin=2 \errorcontextlines10000

7.1. Layout The text width is wide enough for 72-character code lines.

146 \parindent=2em \hsize=36\fontdimen6\tentt \advance\hsize\parindent

A largish left margin is used set—perhaps readers will find it useful for writing comments.

147 \hoffset=210mm \advance\hoffset-\hsize \advance\hoffset-25mm
148 \newdimen\leftmargin \leftmargin=\hoffset \advance\leftmargin-25mm
149 \advance\hoffset-1in

The vertical margins are set so that if the file is formatted for A4 (297 mm) paper and fed into a driver which thinks it is formatted for American Letter (11′′) paper no information will be lost. We allow for 7 mm unprintable area at the edges of paper (this is reasonable for a laser printer), and for the headline being 24 pt + 6 pt − 10 pt = 20 pt above the top of the text. (The 6 pt allows for the height of the headline, which is in 7-pt type.)

150 \voffset=297mm \advance\voffset-11in
151 \advance\voffset7mm \advance\voffset20pt
152 \vsize=297mm \advance\vsize-2\voffset \advance\vsize-2\baselineskip
153 \advance\voffset-1in

We use ragged-right (ranged-left) printing. (Readability studies have indicated that ragged-right printing is, if anything, easier to read than with both sides aligned.) To do this we remove the stretch and shrink components of the font we use, and make the right margin stretchable.

154 \def\do#1{\fontdimen3#1=0pt \fontdimen4#1=0pt }
155 \do\tenrm \do\tenit \do\tenbf
156 \rightskip=0pt plus 2em minus 2pt \adjdemerits=0 \tolerance=500
157 \parfillskip=0pt plus 0.8\hsize

7.2. Title and headline The author and title info are printed at the top of the first page—in lieu of a separate title page—and the title goes in the headline of subsequent pages, followed by the version ID and date as extracted from the RCS identification string. If the file locker is non-empty (i.e., the file is being edited), an ‘e’ is appended to the version ID in the headline.

158 \def\author#1{\def\theauthor{#1}} \author{P. Damian Cugley}
159 \font\titlefont=cmssbx10 scaled \magstep3
160 \def\title#1{
161  \vglue2\baselineskip
162  \begingroup \parindent-\leftmargin \hyphenpenalty10000 \titlefont
163 #1\smallskip
164  \endgroup
165  \leftline{Version \fileversion,
166  last changed \fileauthor\space \mungedate\filedate}
167  \ifx\filelocker\empty\else
168  \leftline{Edited by \filelocker\unskip\ \today}
169  \fi
170  \medskip
171  \hrule \vskip1pt \hrule \vskip1pt \hrule \vskip1pt \hrule
172  \medskip
173  \rightline{\theauthor}
174  \bigskip
175  \headline{\hfil\global\headline{%
176  \sevenbf \hskip-\leftmargin \themarks \hfil #1
177  \fileversion
178  \ifx\filelocker\empty
179  \space\mungedate\filedate
180  \else
181  e\space\today
182  \fi

183 }} 184 }

185 \def\themarks{%
186  \ifx\firstmark\empty\else
187  \firstmark
188  \ifx\firstmark\botmark \else --\botmark \fi
189  \fi

190 }

Some macros for printing the date. The \mungedate macro converts from RCS’s format (using slashes) to ISO format (using short dashes).

191 \def\today{\mungedate{\number\year/\twodigits\month/\twodigits\day}}
192 \def\twodigits#1{\ifnum#1<10 0\fi \number#1}
193 \def\mungedate#1{$ \textfont0=\font \mathcode` /="7B #1 $}

7.3. Sectioning The sectioning commands do no more than generate numbered headings—there are no tables of contents or marks required.

194 \font\secfont=cmssdc10 at 12pt
195 \font\subsecfont=cmssi10
196 \newcount\seccount
197 \newcount\subseccount
198 \def\section#1{%
199  \advance\seccount1 \subseccount=0
200  \bigbreak
201  \begingroup
202  \secfont \hyphenpenalty=10000 \tolerance=1000
203  \rightskip=0pt plus 2em
204  \noindent\llap{\number\seccount.\enspace}#1
205  \medskip
206  \endgroup
207  \noindenttrue

208 }

209 \def\subsec#1{%
210  \advance\subseccount1
211  \medbreak
212  \begingroup
213  \subsecfont \hyphenpenalty=10000 \tolerance=1000
214  \rightskip=0pt plus 2em
215  \noindent\llap{\number\seccount.%
216  \number\subseccount.\enspace}#1
217  \nobreak\smallskip
218  \endgroup
219  \noindenttrue

220 }

Here’s a simple hack to remove the indentation of the first paragraph following a section heading.

221 \newif\ifnoindent
222 \everypar{%
223  \ifnoindent
224  {\setbox0=\lastbox}\global\noindentfalse
225  \fi

226 }

7.4. Footnotes Plain TEX’s footnote command is adjusted to use a new parameter to specify the value of \rightskip in footnotes.* It was the trickiness involved in changing definitions like this that persuaded me to not use ‘@’ characters in macro files.

227 \newskip\footnoterightskip \footnoterightskip=\rightskip

* The default definition sets \rightskip to 0 pt, which is moderately disastrous given that spaces cannot stretch or shrink.

228 \catcode` @=11
229 \def\vfootnote#1{\insert\footins\bgroup
230 \interlinepenalty\interfootnotelinepenalty
231 \splittopskip\ht\strutbox % top baseline for broken footnotes
232 \splitmaxdepth\dp\strutbox \floatingpenalty\@MM
233 \leftskip\z@skip \rightskip\footnoterightskip
234 \spaceskip\z@skip \xspaceskip\z@skip \rm
235 \textindent{#1}\footstrut\futurelet\next\fo@t}
236 \catcode` @=11

This definition is a drop-in replacement for the one in TEXbook Appendix B (p. 363 in my copy).

7.5. Miscellaneous abbreviations

The \< macro for writing the names of syntactic entities (as ⟨foo⟩) is like in TEXbook Appendix E. The \cs macro prints a control sequence name; I use this where the easier dodge of writing |\foo| won’t work (e.g., in the argument to another macro). Similarly \arg places its argument in typewriter-style braces. The \pt macro works in maths mode.

237 \def\<#1>{\leavevmode\hbox{$\langle$\it#1\/$\rangle$}}
238 \def\cs#1{\leavevmode\hbox{\tt\char` \\#1}}
239 \def\arg#1{\leavevmode\hbox{{\tt\char` \{}#1{\tt\char` \}}}}
240 \def\pt{\,{\rm pt}} \def\em{\,{\rm em}} \def\mm{\,{\rm mm}}
241 \def\inch{´ ´ }

To make a version of the LaTEX logo that should work acceptably with a range of fonts (LaTEX, LaTEX, LaTEX), I first define a macro for raising letters until they are flush with the prevailing cap height. I also loosen up the ‘TEX ’ logo (so that it also works with sanserif fonts) and make it set the spacefactor to 1000 (to override TEX’s assumption that a full stop following a capital X isn’t the end of a sentence).

242 \def\flushtop#1{%
243  \leavevmode
244  \begingroup
245  \setbox0=\hbox{#1}\setbox2=\hbox{X}%
246  \dimen0=\ht2 \advance\dimen0-\ht0
247  \raise\dimen0 \box0
248  \endgroup

249 }

250 \def\TeX{T\kern-0.1em \lower0.5ex \hbox{E}X\spacefactor1000 }
251 \def\LaTeX{L\kern-0.1em \flushtop{a}\TeX}

Occasional paragraphs are written in italics and marked with ‘note’.

Note The command says ‘notepar’ not just ‘note’ because in my macro files I use

\note for numbered footnotes ( TEXbook, Exercise 15.12).* Notes may occasionally be more than one paragraph long.

252 \def\notepar{
253  \smallskip
254  \begingroup
255  \leftskip2\parindent \advance\rightskip\parindent minus \parindent
256  \noindent\kern-\parindent{\bf Note}\quad\it
257  \ignorespaces

258 }

259 \def\endnotepar{
260  \smallbreak
261  \endgroup

262 }

* Footnotes in dtx files are produced with the plain \footnote command.

The file is finished with an \endinput command; this can be used by other prgorams to check that this file hasn’t been truncated by TEX coming to a halt part-way through processing the dtx file.

 263 \endinput